Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 53     Entry time: 17 Apr 2025, 01:39
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 18.8 keV, 260 uJ pulse energy

Optical laser delivery

  • No optical laser

Achievements/Observations

  • Experimental setup optimized, both instrumental side and PIXIE side
  • X-ray beam focused to the PIXIE and transmission measurement was done with several different PIXIE operation conditions.

Issues

  • No beam for ~4.5 hours due to ZZ.
  • X-ray beam not found with IRDa InlineMic after focusing X-ray beam to IRDa sample plane.
  • InlineMic Y motor is in error states. It has to be fixed during the day.
  • upstream kapton window for the downstream vacuum pipe broken. The glue has become loose. No damage by X-ray was visible.
ELOG V3.1.4-7c3fd00